OpenStudy (anonymous):

Please solve and explain. Need to find slope and solve it through elimination and substitution. Two brothers are building bridges. Diego started building his bridge that rises 7m up and runs at 4m going east from one side of the river. Daniel build his bridge on the opposite side that rises 3m up and runs at 4m going west.

OpenStudy (anonymous):

Whats the question?

OpenStudy (anonymous):

Use pythagorus? a^2+b^2=c^2

OpenStudy (anonymous):

Uhm, the questions are: a) What is the slope? b) Find x and y through elimination and substitution.

OpenStudy (anonymous):

And please explain how you did it. :< I'm not good in Math.

OpenStudy (anonymous):

Ah, Zhang. I'll use pythagorus for finding the slope? Or for finding the x and y?

OpenStudy (dumbcow):

slope = rise/run Diego: assuming east means positive direction slope = 7/4 Daniel: slope = 3/-4

OpenStudy (anonymous):

Ah! Thanks, Cow. owo